# Survey of Load-Altering Attacks Against Power Grids: Attack Impact, Detection and Mitigation

### v1

@@ IV-B Data-Driven Approaches
With the rapid advancements in artificial intelligence-based techniques in recent years, several studies have adopted ML-based methods to detect and localize LAAs. As LAA affects the load profile of the system, an intuitive method is to monitor and analyze the load data recorded by smart meters. [91] analyzed this data and showed that compromised loads can be detected by performing a fast Fourier Transform (FFT) and spectral analysis of the signals. A windowed FFT (W-FFT) was then used to detect an attack. The authors also investigated the feasibility of using cross-correlation (CC) and windowed cross-correlation (W-CC) to analyze signals. They showed that utilizing both time-domain and frequency-domain signals helps improve the performance of LAA detections.
A time-delay neural network was designed in [92], whose inputs are the power demand profiles. Diverse load profiles were tested, and results showed that the detector could precisely recognize even LAAs that activate only a small number of electric water heaters. [93] pre-processed the load demand data into a feeder loading abnormal power spectrum index. This index was then fed into a long-short-term memory (LSTM) neural network to detect the anomalies. Another data-driven method was proposed in [94] to detect LAAs in distribution networks by exploiting the spatial and temporal properties of the load data. A graph attention network (GAT) and LSTM were employed to capture the spatial correlation and temporal correlations, respectively. The proposed model outperformed existing methods in detecting and localizing LAAs, demonstrating robustness against noise and outliers, as well as scalability for large-scale systems.
It may not always be feasible to measure every real-time load profile. Existing work suggests that the sampling rate of meters could be exploited to launch stealthy load attacks. For instance, if attackers know the meter’s sampling rate, they can synchronize the switching of power circuits to occur between meter readings, leading to inaccurate power consumption calculations [98]. An alternative is to leverage the grid frequency and phase angle fluctuations monitored by PMUs installed to detect and localize LAAs. Based on this, [95] proposed the use of a 2-D convolutional neural network (CNN) classifier and reconstruction decoder to detect and localize DLAAs. The proposed method demonstrated good performance, even in the presence of noise or data loss in the measurements. Focusing on the LAAs utilizing EVCS, an unsupervised deep learning-based mechanism using CNN and EV charging profiles was proposed in [84] to detect malicious attacks.
@@ IV-C Hybrid Approaches
Given the shortcomings of both data-driven and model-based methods, recent work has explored hybrid approaches that combine the strengths of both. These approaches integrate the automatic fitting capabilities of data-driven methods with the physical insights provided by model-based techniques, offering a more balanced and effective solution.
Inspired by the model-based approaches, authors in [96] followed the idea of designing observers to estimate the state of the system, but in the frequency domain. The power system dynamics under attacks were represented using a state-space model, which was then transformed into the frequency domain. An optimization problem constrained by the frequency domain system model was then formulated in order to localize the attacks. The authors also pointed out the feasibility of applying the approach in wide area monitoring systems (WAMS), by implementing it in a hierarchical fashion.
@@ V Attack Mitigation
Following detection, the destabilizing effects of LAAs must be mitigated. Attack mitigation schemes can be either preventive or reactive. Preventive methods aim at deploying security reinforcements offline such that the attacker cannot launch destabilizing LAAs. For instance, reinforcing security features of load devices (e.g., adding hardware or software security features) can prevent load compromises and thus limit the attacker’s ability to launch destabilizing LAAs. However, deploying these security measures may incur significant economic costs upfront. Thus, the second category of reactive methods aims to develop online mitigating measures that will enhance the ability of power grids to withstand the adverse impacts of LAAs. This could, for instance, include generator dispatch from fast-acting resources to correct the imbalance caused by LAAs. Third, hybrid approaches include a combination of offline deployments and online operation of these devices to counter the adverse effects of LAAs. Table V summarizes the existing mitigation methods.
TABLE V: Classification of mitigation methods.
Ref Type grid Defensive tool
[54] Preventive Transmission Securing loads
[102] Preventive Transmission Robust operating points
[103] Preventive Transmission Battery
[104] Preventive Transmission Reactive power compensation
[105] Preventive Distribution Edge computing infrastructure
[106] Reactive Transmission EV charge/discharge
[107] Reactive Transmission Frequency droop control
[108] Reactive Transmission Load shedding
[48] Reactive Distribution Reconfiguration
[109] Reactive Transmission Communication links
[38] Hybrid Transmission Secondary frequency control
[67] Hybrid Distribution Soft open points
[110] Hybrid Transmission Securing devices & Optimal load management
@@ V-A Preventive Mitigation
To minimize successful LAAs, [105] proposed an edge-computing infrastructure that dynamically enforces permissions for each IoT device (e.g., read and execute commands), preventing unauthorized control commands on critical devices and minimizing disruption. Edge servers near IoT devices reduce latency and enable quick responses to attacks.
@@ V-B Reactive Mitigation
The flexibility provided by EVs is proposed to mitigate LAAs in [106]. The EV control actions (such as charge/discharge actions) to compensate for the LAA imbalance are optimized by using $H-2$ and $H-\infty$-based robust control techniques. The work also takes the uncertainties like EV users’ behavior into account. The proposed scheme is shown to keep the system’s frequency in the stable range, thus preventing subsequent contingencies such as line tripping or generator relay tripping.
@@ V-C Hybrid Mitigation
In [110], the work explored a game-theoretic defense that integrates both the cyber and physical aspects of the power grid. At the cyber layer, the attacker aims to spread the malware in the IoT devices and the defender aims to secure the devices (e.g., by software/firmware updates etc.). At the physical layer, the attacker aims to launch LAAs that maximize the impact of the attack, while the defender counteracts it by optimal load management and control strategies to maintain the stability of the grid. Results show that the formulation enables the operator to maintain the system’s normal operation under the strategic botnet attack and improve the grid’s integrative resiliency at both cyber and physical layers.
@@ VI Conclusions & Future Research
In this work, we provide a comprehensive overview of the LAAs against power grids. These attacks can lead to significant disruptions, threatening the stability of the grid, as well as creating opportunities for attackers to manipulate energy market prices for financial gain. Given these risks, it is essential to develop advanced and comprehensive security frameworks. Research is focusing on improving LAA detection and localization techniques, leveraging both model-based and data-driven approaches to enhance real-time monitoring capabilities. Hybrid methods that integrate ML with physics-based models offer promising potential for detecting attacks while maintaining accuracy and speed. Additionally, there is a pressing need to explore effective mitigation strategies that can be deployed both preventively and reactively. These strategies include robust cybersecurity protocols, the use of energy storage for stabilization, network reconfiguration, and strategic coordination between DERs and IoT-enabled devices.
Several open research problems remain, offering fertile ground for future exploration. These challenges span across multiple dimensions, including the need for integrated cyber-physical risk analyses, the economic implications of LAAs on energy markets, and the unique vulnerabilities introduced by RES. Additionally, future work must focus on developing more resilient recovery strategies and enhancing detection and mitigation frameworks, leveraging advanced techniques such as game theory and ML.
@@
(1) Joint Cyber-Physical Risk Analysis: Although several works have analyzed the risks associated with LAAs, there remains a gap in understanding the joint cyber-physical impacts, particularly in the context of DLAAs. Existing research has primarily focused on either cyber or physical aspects, but an integrated analysis is crucial. For example, [110] proposed a two-stage framework where they first analyzed cyber-risks using an epidemic model to track botnet propagation, followed by an examination of the physical impacts on the power grid. This comprehensive approach, which combines game-theoretic formulations to model both the attacker’s strategy and the physical destabilization of the grid, is a promising avenue for future research. Future work should explore similar integrated frameworks to assess cyber-physical risks and optimize joint defense strategies, especially in systems with high penetration of IoT devices and DERs.
@@
(2) LAA Market Manipulation and Economic Impacts: LAAs present significant threats to energy markets by enabling attackers to manipulate market prices through changes in power demand. These attacks can lead to extreme volatility, with prices surging or plummeting based on the attacker’s intent. Such market disruptions not only destabilize grid operations but can also lead to financial losses for market participants. While the feasibility of manipulating energy markets through LAAs has been demonstrated [80], the economic impacts remain an underexplored area. Future research should focus on understanding how LAAs affect electricity markets and on designing resilient market frameworks that can detect and mitigate such manipulations. Developing real-time market response mechanisms and security protocols to ensure market integrity during and after LAAs is another key direction.
@@
(3) LAAs in Renewable Energy-Rich Systems: The integration of RES introduces unique challenges for grid stability, particularly in the face of LAAs. The inherent variability and intermittency of RES make the power grid more susceptible to LAAs. For instance, an attacker could exploit periods of low RES generation by launching an LAA [24], triggering an increase in demand that the grid cannot meet due to insufficient backup from conventional generators. Similarly, during periods of high RES output, LAAs could result in overproduction, leading to instability in grid frequency and voltage. Future research should focus on how LAAs interact with the variability of RES and on developing controls to mitigate these risks. Integrating advanced forecasting techniques and real-time monitoring could help grid operators maintain stability in both normal and attack scenarios.
@@
(5) Enhanced Detection and Mitigation via Game Theory and ML: Given the evolving sophistication of LAA strategies, future research should focus on developing more robust detection and mitigation mechanisms. Game-theoretic approaches, as discussed in [112, 48, 113], can model the interaction between attackers and defenders, enabling the prediction of optimal attack and defense strategies. Additionally, combining game theory with ML can improve the accuracy of detection and help anticipate future attacks. Hybrid detection methods combining ML to identify abnormal power demand patterns with physics-based models to maintain system stability present a promising approach for enhancing LAA defense strategies.
By addressing these research gaps, the security and resilience of future power grids can be significantly enhanced, ensuring both grid stability and market integrity.
@@ Acknowledgments
This work has been supported in part by PhD Cofund WALL-EE project between the University of Warwick, UK and CY Cergy Paris University, France, EPSRC under Grant EP/S035362/1, and King Abdullah University of Science and Technology (KAUST) under Awards No. ORFS-2022-CRG11-5021, No. RFS-OFP2023-5505, and the KAUST Center of Excellence for Renewable Energy and Storage Technologies (CREST), under award number 5937.

### v2

@@ IV-B Data-Driven Approaches
With the rapid advancements in artificial intelligence-based techniques in recent years, several studies have adopted ML-based methods to detect and localize LAAs. As LAA affects the load profile of the system, an intuitive method is to monitor and analyze the load data recorded by smart meters. [97] analyzed this data and showed that compromised loads can be detected by performing a fast Fourier Transform (FFT) and spectral analysis of the signals. A windowed FFT (W-FFT) and windowed cross-correlation (W-CC) were then used to detect an attack. They showed that utilizing time and frequency-domain signals could improve the performance of LAA detections.
A time-delay neural network was designed in [98], whose inputs are the power demand profiles. Results showed that the detector could precisely recognize even LAAs that activate only a small number of electric water heaters. [99] pre-processed the load demand data into a feeder loading abnormal power spectrum index. This index was then fed into a long-short-term memory (LSTM) network to detect the anomalies. Another method was proposed in [100] to detect LAAs in distribution networks by exploiting the spatial and temporal properties of the load data. A graph attention network (GAT) and LSTM were employed to capture the spatial and temporal correlations, respectively. The proposed model outperformed existing methods in detecting and localizing LAAs, demonstrating robustness against noise and outliers.
It may not always be feasible to measure every real-time load profile. Existing work suggests that the sampling rate of meters could be exploited to launch stealthy load attacks. For instance, if attackers know the meter’s sampling rate, they can synchronize the switching of power circuits to occur between meter readings, leading to inaccurate power consumption calculations [104]. An alternative is to leverage the grid frequency and phase angle fluctuations monitored by PMUs installed to detect and localize LAAs. Based on this, [101] proposed the use of a convolutional neural network (CNN) classifier and reconstruction decoder to detect and localize DLAAs. The proposed method demonstrated good performance, even in the presence of noise or data loss in the measurements. Focusing on the LAAs utilizing EVCS, an unsupervised deep learning-based mechanism using CNN and EV charging profiles was proposed in [58] to detect malicious attacks.
@@ IV-C Hybrid Approaches
Figure 3: Comparison among model-based, data-driven and hybrid approaches.
Given the shortcomings of both data-driven and model-based methods, recent work has explored hybrid approaches that combine the strengths of both, as illustrated in Fig. 3. These approaches integrate the automatic fitting capabilities of data-driven methods with the physical insights provided by model-based techniques, offering more balanced and effective solutions.
Inspired by the model-based approaches, [102] followed the idea of designing observers to estimate the state of the system. The power system dynamics under attacks were represented using a state-space model, which was then transformed into the frequency domain. An optimization problem constrained by the frequency domain system model was then formulated in order to localize the attacks. The authors also pointed out the feasibility of applying the approach in wide area monitoring systems (WAMS), by implementing it in a hierarchical fashion.
@@ V Attack Mitigation
Once detected, the destabilizing effects of LAAs need be addressed. Attack mitigation schemes can be either preventive or reactive. Preventive methods aim at deploying security reinforcements offline such that the attackers cannot launch destabilizing LAAs. For instance, reinforcing security features of load devices (e.g., adding hardware or software security features) can prevent load compromises and thus limit the attacker’s ability to launch destabilizing LAAs. However, deploying these security measures may incur significant economic costs upfront. Thus, the second category of reactive methods aims to develop online mitigating measures that will enhance the ability of power grids to withstand the adverse impacts of LAAs. This could, for instance, include generator dispatch from fast-acting resources to correct the imbalance caused by LAAs. Third, hybrid approaches include a combination of offline deployments and online operation of these devices to counter the adverse effects of LAAs. Table VI summarizes the existing mitigation methods.
TABLE VI: Classification of mitigation methods.
Type Grid Ref Defensive tool
Preventive Transmission [60] Securing loads
[75] Increasing data privacy, Sharing altered data, and installing non-intrusive load monitoring algorithm
[108] Robust operating points
[109] Battery
[110] Reactive power compensation
Distribution [111] Edge computing infrastructure
[112] ML-based market clearing
[113] Securing critical aggregators
Reactive Transmission [114] EV charge/discharge
[115] Frequency droop control
[116] Load shedding
[117] Communication links
[118] Canceling selected bids
Distribution [57] Reconfiguration
[119] Fraud data detection
Hybrid Transmission [49] Secondary frequency control
[120] Securing devices & Optimal load management
Distribution [76] Soft open points
@@ V-A Preventive Mitigation
To reduce the risk of successful LAAs, [111] proposed an edge-computing infrastructure that dynamically enforces permissions for each IoT device (e.g., read and execute commands), preventing unauthorized control commands on critical devices and minimizing disruption. Edge servers near IoT devices reduce latency and enable quick responses to attacks.
While most existing studies focus on mitigating the destabilizing effects of LAAs, a few works focus on mitigating LAA impacts on the electricity market or market-based mitigation. [112] uses an LSTM-based method to capture temporal patterns the energy prices based on the import and export prices set by retailers and communicated to households. Once the LTSM is trained, a confidence interval is established based on prediction errors, and any price prediction that falls outside this range is flagged as a potential attack. [75] provides information on the most effective countermeasures against market manipulation attacks, including sharing detailed market data only with market players such as independent system operators, releasing redacted or altered versions of market data or even delaying the release of full data sets to the public. These methods limit access to historical market data and enhance data privacy. As a result, implementing real-time attacks against the market becomes difficult. Furthermore, installing non-intrusive load monitoring (NILM) at the household level can be an effective countermeasure. For example, NILM can detect the unusual operation of home appliances during the periods when the homeowner is not present. [113] proposes a mixed-integer nonlinear programming optimization problem to design an attack against the electricity market that causes the highest tariff increment and congestion issues in the network. Based on this formulation, they locate the aggregators from which the attack can be launched with the least effort and propose securing these aggregators to mitigate the attacks.
@@ V-B Reactive Mitigation
The flexibility provided by EVs is proposed to mitigate LAAs in [114]. The EV control actions (such as charge/discharge actions) to compensate for the LAA imbalance are optimized using $H-2$ and $H-\infty$-based robust control techniques. The work also takes the uncertainties like EV users’ behavior into account. The proposed scheme keeps the system’s frequency in the stable range, thus preventing subsequent contingencies such as line tripping or generator relay tripping.
@@
To counteract the attacks on electricity markets, [119] propose a scheme to detect and penalize agents that manipulate the electricity market by introducing a fraud detection tool that compares real-time bids with historical data. Any agent (node) identified as an adversary will incur a penalty. Furthermore, [118] considers that the operator has historical data of the bids submitted by the market players. Based on this when there is an attack, they can estimate the number of devices under attack as well as the impact of the attack. To compensate for the impact of the attack on electricity price (which is the increment in the price), the operator drops some of the highest-priced bids to correct the effect of the attack.
@@ V-C Hybrid Mitigation
In [120], the work explored a game-theoretic defense that integrates both the cyber and physical aspects of the power grid. At the cyber layer, the attacker seeks to spread malware in IoT devices, while the defender focuses on securing them (e.g., through software and firmware updates). At the physical layer, the attacker aims to launch LAAs that maximize the impact of the attack, while the defender counteracts it by optimal load management and control strategies to maintain the stability of the grid. Results show that the formulation enables the operator to maintain the system’s normal operation under the strategic botnet attack and improve the grid’s integrative resiliency at both cyber and physical layers.
@@ VI Conclusions & Future Research
LAAs pose a significant cybersecurity threat to modern power grids, exploiting IoT-enabled high-wattage consumer devices to manipulate power consumption at scale. This paper provides a comprehensive introduction to LAAs by classifying them based on various evaluation criteria, offering a deeper understanding of their forms and impact. The survey also systematically investigates the impacts of LAAs on power system operations, revealing the primary consequences: grid frequency and voltage instability, and electricity market disequilibrium. A comparative analysis of detection methodologies is conducted, encompassing model-based analytical techniques, data-driven machine-learning approaches, and hybrid detection frameworks. The mitigation framework is comprehensively evaluated through two complementary dimensions: 1) preventive security measures, including cryptographic protection mechanisms for IoT-enabled smart loads, and 2) reactive countermeasures incorporating adaptive demand-side management protocols and real-time emergency load shedding algorithms.
The evolving landscape of LAAs presents significant cybersecurity risks to power systems, necessitating proactive policy interventions, regulatory frameworks, and industry practices. The revealed vulnerability of IoT-enabled high-wattage devices underscores the urgency for implementation of robust security measures, such as IEEE 2030.5 (Smart Energy Profile 2.0) [122] security extensions, which provides secure communication mechanisms for DERs. Moreover, recent governmental responses highlight the importance of securing smart appliances and demand-side response mechanisms to prevent large-scale cyber threats. For instance, the Cyber Resilience Act by the European Commission aims to enhance the security of digital products by enforcing stricter cybersecurity requirements across the supply chain [123]. Similarly, in the U.S., the NIST cybersecurity labeling program [124] focuses on improving consumer awareness and incentivizing manufacturers to implement stronger security measures in consumer devices, contributing to the overall resilience of energy systems. UK’s smart energy security framework also emphasizes securing demand-side response technologies and ensuring compliance with cybersecurity standards for energy smart appliances [125].
To effectively mitigate LAA risks, it is imperative that policymakers establish enforceable cybersecurity baselines that integrate LAA-specific threat models into existing grid security frameworks. Ensuring compliance with standardized security protocols, such as IEEE 2030.5, plays a crucial role in enhancing interoperability and protecting IoT-enabled loads from large-scale coordinated LAAs. Additionally, industry stakeholders should prioritize the deployment of advanced anomaly detection mechanisms and adaptive defense strategies to counter evolving threats. A multidisciplinary approach that combines regulatory enforcement, technological innovation, and cross-sector collaboration is essential to enhancing grid resilience against cyber-physical threats. By integrating these measures, power systems can transition toward a more secure, intelligent, and adaptive operational paradigm, effectively mitigating the impact of LAAs while maintaining system stability and reliability.
Several open research problems remain, offering fertile ground for future exploration.
@@
(1) Joint Cyber-Physical Risk Analysis: Although several works have analyzed the risks associated with LAAs, there remains a gap in understanding the joint cyber-physical impacts, particularly in the context of DLAAs. Existing research has primarily focused on either cyber or physical aspects, but an integrated analysis is crucial. For example, [120] proposed a two-stage framework where they first analyzed cyber-risks using an epidemic model to track botnet propagation, followed by an examination of the physical impacts on the power grid. Future work should assess cyber-physical risks and develop optimized joint defense strategies, particularly in systems with high IoT and DER penetration.
@@
(2) LAA Market Manipulation and Economic Impacts: LAAs present significant threats to energy markets by enabling attackers to manipulate market prices through changes in power demand. These attacks can lead to extreme volatility, with prices surging or plummeting based on the attacker’s intent. While the feasibility of manipulating energy markets through LAAs has been demonstrated [73], the economic impacts, real-time market response mechanisms, and security protocols to ensure market integrity during and after LAAs, remain underexplored areas.
@@
(3) LAAs in Renewable Energy-Rich Systems: The inherent variability and intermittency of RES make the power grid more susceptible to LAAs. For instance, an attacker could exploit periods of low RES generation by launching an LAA [27], triggering an increase in demand that the grid cannot meet due to insufficient backup from conventional generators. Similarly, during periods of high RES output, LAAs could result in overproduction, leading to instability in grid frequency and voltage. Future research should focus on integrating advanced forecasting techniques and real-time monitoring which could help grid operators maintain stability during LAAs.
@@
(5) Enhanced Detection and Mitigation via Game Theory and ML: Given the evolving sophistication of LAA strategies, future research should focus on developing more robust detection and mitigation mechanisms. Game-theoretic approaches, as in [126, 57, 127], can model the interaction between attackers and defenders, enabling the prediction of optimal attacks and defenses. Additionally, combining game theory with ML to identify abnormal power demand patterns with physics-based models to maintain system stability presents a promising approach for enhancing LAA defense strategies.
@@ Acknowledgments
This work has been supported in part by PhD Cofund WALL-EE project between the University of Warwick, UK and CY Cergy Paris University, France, EPSRC under Grant EP/S035362/1, and King Abdullah University of Science and Technology (KAUST) under Awards No. ORFS-2022-CRG11-5021 and No. RFS-OFP2023-5505.